OpenStudy (anonymous):

A triangle has a base x and height y. The hypotenuse is 12 and the area is 24. What is x and y using quadratic equation to solve?

OpenStudy (raden):

there are 2 equations : x^2 + y^2 = 12^2 1/2 * x * y = 24

OpenStudy (raden):

from the 2nd : get 1/2 x y = 24 xy = 48 y = 48/x

OpenStudy (raden):

subtitute y = 48/x into the 1st equation : x^2 + y^2 = 12^2 x^2 + (48/x)^2 = 144 x^2 + 2304/x^2 = 144 then let u = x^2, so the equation can be u + 2304/u = 144 multiply by u on both sides, get u^2 + 2304 = 144u or u^2 - 144u + 2304 = 0 solve for u, first. can you ?

OpenStudy (gorv):

by pythagorse theorm base^2+height^2=hypotenuse^2